# Deciphering Functions of Neurons in Vision-Language Models

### v1

@@ 4.1 Neuron activation dataset
Figure 6: Distribution of the four types of neurons when we treat the determination of the neuron type as a classification problem.
@@ 4.3 Distributions of different types of neurons
We show the histogram for the probability of $p_{v}$, $p_{t}$, $p_{m}$, and $p_{u}$ over all the neurons, respectively in Figure 5(a)-(d).
In Figure 5 (a), we can see that there is a small portion of neurons having very high probability of being visual neurons (where $p_{v}$ approaches 1). There are about 5.4% neurons having probability higher than 80% to be visual neurons. From Figure 5 (b)-(d), there are about 14.2%, 5.5%, and 5.7% neurons having a probability higher than 80% to be text neurons, multi-modal neurons, and unknown neurons, respectively.
When we consider the identification of neuron type as a classification problem (by recognizing the type of high probability as the neuron type), we obtain the class distribution for the four types of neurons from all the neurons (352,226) as shown in Figure 6. We can see that text neurons have the highest frequency while the other three types of neurons exhibit similar proportions among all the neurons. In fact, as shown in Figure 5 (a), besides a small portion of neurons that can be clearly identified of their types, there are abundant neurons that are less confident to be clearly identified as specific neurons. For example, a neuron may have the probability of $p_{v}=0.4$, $p_{t}=0.1$, $p_{m}=0.3$, and $p_{u}=0.2$ of being the four types.
@@ 4.7 Impact of pruning unknown neurons
We investigate the impact of pruning unknown neurons on LVM by evaluating its performance on visual question answering (VQA) task. The results show that dropping unknown neurons brings smaller performance drop than randomly dropping neurons (see the Appendix B.1).
@@ C.2 Distributions of different types of neurons
When we consider the identification of neuron type as a classification problem (by recognizing the type of high probability as the neuron type), we obtain the class distribution for the four types of neurons from all the neurons (458,752) as shown in Figure 22. In comparison to LLaVA 1.5, the InternVL 2.5 exhibits a notable increment in the quantity of visual neurons and a significant decrease in the proportion of text neurons. This phenomenon may be attributed to the enhanced emphasis placed by InternVL 2.5 on bolstering its visual data processing capabilities (Chen et al., 2024), particularly with respect to the handling of multi-image and video inputs. During the full model instruction tuning phase (stage 2), the entire model is optimized. In fact, as shown in Figure 21 (a), besides a small portion of neurons that can be clearly identified of their types, there are abundant neurons that are less confident to be clearly identified as specific neurons.
@@ C.3 Neuron distributions across layers
Figure 22: Distribution of the four types of neurons of InternVL 2.5. We treat the determination of the neuron type as a classification problem.
@@
We can see that for visual neurons and text neurons, they present high frequency in the early and middle layers while low frequency in the high layers. In contrast, the multi-modal neurons have higher frequency in the high layers. That may because in the early and middle layers, more visual neurons and text neurons are learned to focus on processing the individual modality. In the high layers, more multi-modal neurons are learned to jointly deal with visual and text information, enabling efficient reasoning of visual and text information. Interestingly, this distribution trend is similar to that of LLaVA 1.5.

### v3

@@ 4.1 Neuron activation dataset
Figure 6: Distribution of the four types of neurons (visual-prone neurons, text-prone neurons, multimodal-prone neurons, unknown neurons) when we treat the determination of the neuron type as a classification problem.
@@ 4.3 Distributions of different types of neurons
We show the histogram for the probability of $p_{v}$, $p_{t}$, $p_{m}$, and $p_{u}$ over all the neurons, respectively in Figure 5(a)-(d). Figure 5 (a) shows that there is a small portion of neurons having very high probability of being visual neurons (where $p_{v}$ approaches 1). There are about 5.4% neurons having probability higher than 80% to be visual neurons. From Figure 5 (b)-(d), there are about 14.2%, 5.5%, and 5.7% neurons having a probability higher than 80% to be text neurons, multi-modal neurons, and unknown neurons, respectively.
When we consider the identification of neuron type as a classification problem by recognizing the type of high probability as the neuron type, we obtain the class distribution for the four types of neurons (we refer to them as visual-prone neurons, text-prone neurons, multimodal-prone neurons, unknown neurons) from all neurons (352,226) as shown in Figure 6. In fact, as shown in Figure 5 (a), besides a small portion of neurons that can be clearly identified of their types, there are abundant neurons that are less confident to be clearly identified as specific neurons. For example, a neuron may have the probability of $p_{v}=0.4$, $p_{t}=0.1$, $p_{m}=0.3$, and $p_{u}=0.2$ of being the four types. Therefore, more precisely, we refer to it as a text-prone neuron. From Figure 6, we can see that text-prone neurons have the highest frequency, while the other three types of neurons exhibit similar proportions among all the neurons.
@@ 4.7 Impact of pruning unknown neurons
We investigate the impact of pruning unknown neurons on LVM by evaluating its performance on visual question answering (VQA) task. The results show that dropping unknown neurons brings smaller performance drop than randomly dropping neurons (see Appendix B.1).
@@ C.2 Distributions of different types of neurons
When we consider the identification of neuron type as a classification problem by recognizing the type of high probability as the neuron type, we obtain the class distribution for the four types of neurons (we refer to them as visual-prone neurons, text-prone neurons, multimodal-prone neurons, unknown neurons) from all neurons (458,752) as shown in Figure 22. In comparison to LLaVA 1.5, the InternVL 2.5 exhibits a notable increment in the quantity of visual-prone neurons and a significant decrease in the proportion of text-prone neurons. This phenomenon may be attributed to the enhanced emphasis placed by InternVL 2.5 on bolstering its visual data processing capabilities (Chen et al., 2024), particularly with respect to the handling of multi-image and video inputs. During the full model instruction tuning phase (stage 2), the entire model is optimized. In fact, as shown in Figure 21 (a), besides a small portion of neurons that can be clearly identified of their types, there are abundant neurons that are less confident to be clearly identified as specific neurons.
@@ C.3 Neuron distributions across layers
Figure 22: Distribution of the four types of neurons (visual-prone neurons, text-prone neurons, multimodal-prone neurons, unknown neurons) of InternVL 2.5. We treat the determination of the neuron type as a classification problem.
@@
For visual neurons and text neurons, they present high frequency in the early and middle layers while low frequency in the high layers. In contrast, the multi-modal neurons have higher frequency in the high layers. That may because in the early and middle layers, more visual neurons and text neurons are learned to focus on processing the individual modality. In the high layers, more multi-modal neurons are learned to jointly deal with visual and text information, enabling efficient reasoning of visual and text information. Interestingly, this distribution trend is similar to that of LLaVA 1.5.